# Rigorous Modeling of Droplet Impact Dynamics on Micro-Structured Surfaces: Nonlocal Theory and SPH Simulation of Pancake Bouncing

### v1

@@ Abstract
The accurate mathematical modeling of droplet impact dynamics on micro-structured surfaces is fundamental to understanding and predicting complex fluid behaviors relevant to a wide range of engineering and scientific applications. In particular, the pancake bouncing phenomenon—systematically studied by Liu et al. (Nature Physics, 2014)—on superhydrophobic micro-structured substrates presents significant theoretical challenges. Central to these challenges is the need to construct rigorous mathematical models that capture the intricate influence of substrate micro/nanostructures on droplet dynamics. This requires the development of robust formulations for surface tension, contact line dynamics, and the interaction forces between fluid and solid structures. In this work, we formulate a nonlocal mathematical framework for the simulation of 3D pancake bouncing on superhydrophobic micro-cone arrays. The model incorporates intermolecular attractive forces to represent droplet surface tension, and we provide a strict theoretical derivation linking these forces quantitatively to the macroscopic surface tension coefficient, thereby circumventing the reliance on empirical parameter tuning. The complex geometry of micro-cone arrays introduces fundamental difficulties in defining local normal directions for contact algorithms. To overcome this, we develop a nonlocal contact repulsion force model that governs fluid-solid interactions and ensures numerical stability under high Weber number conditions. Based on this mathematical foundation, we implement the model using smoothed particle hydrodynamics (SPH), enabling high-precision 3D simulations. Computational experiments, validated against empirical data, confirm the model’s accuracy and robustness, while underscoring the key role of numerical simulation in elucidating droplet-microstructure interactions.
keywords: Pancake bouncing; SPH method; attractive forces; nonlocal contact repulsion
@@ 1 Introduction
Understanding and controlling droplet rebound dynamics is essential for advancing liquid manipulation technologies in various engineering applications. This phenomenon is particularly critical in applications such as combustion systems [1], inkjet printing [2], anti-icing surfaces [3], corrosion-resistant coatings [4] and self-cleaning materials [5]. Recent studies reveal that these technologies have broad applications across fields including aerospace, biomedicine, automotive engineering, and smart textiles [6]. While hydrophobic coatings facilitate droplet removal, achieving directional control and efficient shedding requires not only low surface energy but also carefully designed nanostructures with tailored surface chemistry [7]. At such nano-enhanced interfaces, pancake bouncing [8] emerges as a remarkable phenomenon: droplets impacting nanostructured superhydrophobic surfaces lift off immediately in a flattened state without retracting, reducing contact time by approximately 75% compared to conventional rebound. Precise liquid control is crucial for advanced functional surfaces and devices, necessitating deeper research into droplet rebound dynamics, especially given their significant implications in healthcare and electronics.
@@
The dynamics of droplet rebound has garnered considerable scientific attention, driving a steady stream of key discoveries in interfacial science[9]. Early studies by [10] revealed that droplets on superhydrophobic surfaces in low-pressure environments can self-remove through spontaneous levitation and trampoline-like bouncing. Later, [11] later identified a superhydrophobic-like bouncing regime on thin liquid films, where the rebound dynamics were independent of the underlying substrate. Subsequent work by [12] introduced a facile method to fabricate hierarchical multilayered superhydrophobic surfaces for efficient droplet shedding. [13] further explored symmetry-breaking impacts, showing that ellipsoidal droplets reduce contact time and suppress rebound magnitude. Recent advances have significantly enhanced control over droplet rebound behavior through novel interfacial strategies. [14] proposed a ’golden section’ design criterion to regulate rebound via structural spacing, challenging conventional approaches. [15] developed a leaf-inspired superhydrophobic cantilever for directional bouncing, while [16] reduced contact time by coating droplets with hydrophobic particles. Additionally, through the utilization of innovative tools, a deeper understanding of the underlying mechanisms has been achieved. In their study, [17] utilized high-speed infrared thermography and a three-dimensional transient heat conduction COMSOL model to accurately map the dynamic heat flux distribution during droplet impact on a cold superhydrophobic surface. [18] introduced a novel micro-controlled droplet generator capable of releasing two equally sized water droplets simultaneously on-demand, enabling the investigation of the behavior of two droplets on a dry substrate. [19] reported that controlled self-transport of bouncing droplets can be achieved with the assistance of ultra-smooth surfaces featuring wedge-shaped grooves. [20] uncovered molecular-scale wetting dynamics through in situ characterization, providing fresh insights into droplet-surface interactions. And [21] conducted a comprehensive numerically assisted analysis based on verifiable assumptions such as quasi-stationarities and small Reynolds/Peclet numbers.
@@
Although experimental investigations provide critical observations, numerical modeling is essential for achieving fundamental physical understanding and enabling technological applications. In fact, despite extensive experimental investigations into droplet impact phenomena, the fundamental physics governing collisions with hydrophobic surfaces remains inadequately understood, as noted in literature [22]. On the contrary, numerical modeling offers unique insights into interfacial phenomena. For example, [23] conducted numerical simulations based on the Cahn-Hilliard-Navier-Stokes model to achieve a complete and quantitative study on the effects of interfacial tension when it becomes sufficiently low. Additionally, [24] simulated buoyancy-driven bubble-wall interactions using an adaptive ALE method, capturing the interaction of a rising bubble with a solid wall. Furthermore, [25] investigated the head-on collision of unequal-size droplets of the same liquid on wetting surfaces using direct numerical simulations at different size ratios. Based on the Onsager variational principle, [20] presented a phenomenological model that reveals how the intrinsic material parameters of soft gels dictate phase-separation dynamics. Moreover, fabricating nanostructured substrates for droplet impact studies incurs significant costs, whereas numerical simulations provide a cost-efficient alternative for parametric studies and experimental design optimization. Within this framework, numerical simulations have become pivotal in impact dynamics research. The creation of high-fidelity computational models capable of capturing intricate interfacial processes is essential for both theoretical advancements and practical applications.
@@
Smoothed Particle Hydrodynamics (SPH) method, as a mesh-free Lagrangian method, offers distinct advantages in simulating droplet rebound phenomena. In fact, droplet rebound typically involves large interfacial deformations and topological changes, presenting inherent difficulties for grid-based simulations. Traditional Eulerian approaches, such as the Lattice Boltzmann method (LBM), rely on multiphase domain discretization and phase-field method to accurately capture fluid interfaces. While LBM excels in modeling complex multi-phase flows, its application to gas-liquid systems with high density ratios demands non-trivial modifications to ensure numerical stability [26, 27, 28]. By contrast, SPH circumvents these limitations through its fully Lagrangian mesh-free formulation, eliminating the need for explicit gaseous-phase discretization altogether [29]. This intrinsic capability renders SPH particularly suited for investigating droplet interactions with hydrophobic surfaces. Other alternative particle-based approaches such as molecular dynamics (MD) and dissipative particle dynamics (DPD) suffer from severe scalability limitations in mesoscale droplet simulations due to their demanding resolution requirements and consequent computational overhead. The SPH method achieves an optimal balance between numerical resolution and computational efficiency. Notably, while SPH has seen extensive practical applications in recent years, its theoretical foundations have been progressively strengthened. Significant contributions include Du et al.’s mathematical analysis from a nonlocal perspective [30, 31] and Sun et al.’s stability analysis through an energy-based framework [32, 33, 34]. These theoretical advancements have collectively reinforced SPH’s position as a robust and well-validated methodology for droplet dynamics research.
@@
Surface tension modeling constitutes a critical component of SPH simulation for droplet dynamics, and numerous significant studies have been conducted in this area. Distinct from phase-field methods that regulate interface motion through energy minimization, SPH implementations necessitate the explicit definition of forces acting on interface particles to control interfacial behavior. In reality, researchers have developed two principal methodological approaches. The first approach is the Continuous Surface Force (CSF) formulation [35], which implements a volumetric force at fluid interfaces through local curvature estimation. While this method maintains physical consistency, it demands careful handling of surface normal vectors and curvature calculations near interface regions [36, 37, 38]. Furthermore, this approach faces challenges in simulating droplet topological changes. The second approach involves the particle-particle interaction force (PIF) model [39, 40], which bypasses the need for normal vector and curvature calculations. This method demonstrates versatility in simulating solid-liquid interactions and solid wettability. Despite its practical effectiveness, the second approach dependence on empirically determined strength of the interparticle attraction [41], lacking strong theoretical foundations, constitutes a major limitation. Notice that surface tension emerges from interparticle attractive forces. Thus, when these attraction strengths are empirically predetermined, the resulting surface tension coefficients become inherently uncontrollable, frequently leading to physical inconsistencies. Therefore, developing a method that preserves theoretical consistency while ensuring conciseness and broad applicability has become critically important.
@@
The contact algorithms is an another challenge for stable and precise SPH simulation, particularly for simulating droplet collisions on intricate arrays of microcones. The inherent complexity of droplet impact dynamics, characterized by multiscale solid-liquid interactions, creates escalating challenges for contact algorithm stability as the Weber number increases [42]. The traditional repulsive force algorithms [36, 37, 43], as a popular contact algorithm, require accurate surface normals $\mathbf{n}$ and vertical distance $d_{p}$ to surface. Specifically, the repulsive force $\mathbf{F}_{i}^{\textrm{rep}}$ on particle $i$ satisfies the following relation:
$\mathbf{F}_{i}^{\textrm{rep}}\propto\frac{2|d_{0}-d_{p}|}{(\Delta t)^{2}}m_{i}\mathbf{n}_{i},$ (1)
@@
where $\Delta t$ is time step, $m_{i}$ is particle mass and $d_{0}$ is contact threshold distance. Although this approach demonstrates satisfactory performance for ideally smooth substrates [9], its extension to realistically rough substrates encounters fundamental limitations. A representative case involves micro-structured hydrophobic surfaces [8], as illustrated in Fig. 1.
@@
In fact, due to the non-smoothness of the substrates, it is hard to determinate the local surface normals $\mathbf{n}_{i}$ and vertical distance $d_{p}$. Besides, the virtual particle method is a widely used boundary treatment technique to address kernel truncation and wall penetration issues. However, when used in isolation, the virtual particle method often fails to reproduce physically accurate repulsive forces during droplet rebound. Consequently, the development of a robust contact algorithm capable of handling intricate microcone geometries is critical for reliable simulation of pancake bouncing dynamics.
@@ 2 Governing equations
$\frac{\mathrm{d}\rho}{\mathrm{d}t}$ $=-\rho\nabla\cdot\mathbf{u},$ (2)
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}$ $=-\nabla p+\mu\nabla^{2}\mathbf{u}+\rho\mathbf{g}+\mathbf{f}_{\mathrm{att}}+\mathbf{F},$ (3)
The governing equations involve the fluid velocity field $\mathbf{u}$, hydrodynamic pressure $p$, and local densities $\rho$. The pressure $p$ is calculated by the following state of equation
$p=\frac{c^{2}\rho_{0}}{\gamma}\left(\left(\frac{\rho}{\rho_{0}}\right)^{\gamma}-1\right).$ (4)
We adopt $\gamma=7$ [44] for the equation of state, with $\rho_{0}$ and $c$ denoting the reference densities and the numerical speed of sound, respectively. $\frac{\mathrm{d}}{\mathrm{d}t}$ represents the material derivative where $\frac{\mathrm{d}}{\mathrm{d}t}\equiv\frac{\partial}{\partial t}+\mathbf{u}\cdot\nabla$. $\mu$ denotes dynamic viscosity. The body forces consist of gravitation $\rho\mathbf{g}$, droplet cohesive attractive force $\mathbf{f}_{\mathrm{att}}$, and contact repulsion $\mathbf{F}$ at the boundary interface.
Accurate characterization of cohesive attractive and repulsive interactions is crucial for reliably simulating droplet rebound dynamics on microcone arrays surfaces. In this section, we present a unified mathematical framework that integrates smoothed kernel functions with nonlocal operators to describe both $\mathbf{f}_{\mathrm{att}}$ and $\mathbf{F}$, enabling robust three-dimensional modeling of bouncing behavior.
@@ 2.1 Modeling of droplet cohesion
It is well-established that the surface tension of water droplets originates from intermolecular forces (hydrogen bonding and van der Waals interactions) between water molecules. While intermolecular interactions are physically straightforward, they impose prohibitive computational costs in macroscopic system simulations. For surface tension modeling at continuum scales, cohesive force models offer a viable alternative. For the simplicity of the mathematical representation, we need to introduce the smoothing kernel function $W(\mathbf{x}-\mathbf{x}^{\prime},h)$, which satisfies the following conditions:
@@
  • Symmetric property: $W(\mathbf{x}-\mathbf{x}^{\prime},h)=W(\mathbf{x}^{\prime}-\mathbf{x},h)$;
[…]
Here, $|\cdot|$ represents the Euclidean norm and $h$ is the influence radius of kernel. $W(\mathbf{x},h)$ affects only a specific support domain.
We now proceed to construct a cohesive forces model capable of generating physically accurate interfacial tension. In our work, each Lagrangian particle represents a finite fluid volume element comprising numerous water molecules. Between adjacent particles, we introduce an artificial cohesive force term $\mathbf{f}(\mathbf{x},\mathbf{y})$ to approximate the net effect of microscopic intermolecular forces. Specifically, we denote the attraction of the particle at $\mathbf{y}$ to the particle at $\mathbf{x}$ as follows
$\mathbf{f}(\mathbf{x},\mathbf{y})=A_{\sigma}\rho(\mathbf{y})\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},{R_{1}}),$ (5)
where $A_{\sigma}$ is a parameter dependent on interfacial tension $\sigma$. The specific form of $A_{\sigma}$ will be given later. $\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},{R_{1}})$ in the same direction as $\mathbf{y}-\mathbf{x}$. Thus, by integrating over $\mathbf{y}$, the total attractive force acting on the particle located at position $\mathbf{x}$ is
$\mathbf{f}_{\mathrm{att}}(\mathbf{x})=\int_{\Omega}\mathbf{f}(\mathbf{x},\mathbf{y})d\mathbf{y}=A_{\sigma}\int_{\Omega}\rho(\mathbf{y})\nabla_{\mathbf{x}}W(\mathbf{x}-\mathbf{y},{R_{1}})d\mathbf{y},$ (6)
@@
$V_{1}$$V_{2}$$\mathbf{x}_{0}$$\mathbf{t}$$R_{1}$$R_{1}$$L$

### v2

# An Effective Model for Droplet Impact Dynamics on Micro-Structured Surfaces: Nonlocal Theory and SPH Simulation of Pancake Bouncing

@@ Abstract
The accurate mathematical modeling of droplet impact dynamics on micro-structured surfaces is fundamental to understanding and predicting complex fluid behaviors relevant to a wide range of engineering and scientific applications. In particular, the pancake bouncing phenomenon—systematically studied by Liu et al. (Nature Physics, 2014)—on superhydrophobic micro-structured substrates presents significant theoretical challenges. Central to these challenges is the need to construct effective mathematical models that capture the intricate influence of substrate micro/nanostructures on droplet dynamics. This requires the development of robust formulations for surface tension, contact line dynamics, and the interaction forces between fluid and solid structures. In this work, we formulate a nonlocal mathematical framework for the simulation of 3D pancake bouncing on superhydrophobic micro-cone arrays. The model incorporates intermolecular attractive forces to represent droplet surface tension, and we provide a strict theoretical derivation linking these forces quantitatively to the macroscopic surface tension coefficient, thereby circumventing the reliance on empirical parameter tuning. The complex geometry of micro-cone arrays introduces fundamental difficulties in defining local normal directions for contact algorithms. To overcome this, we develop a nonlocal contact repulsion force model that governs fluid-solid interactions and ensures numerical stability under high Weber number conditions. Based on this mathematical foundation, we implement the model using smoothed particle hydrodynamics (SPH), enabling high-precision 3D simulations. Computational experiments, validated against empirical data, confirm the model’s accuracy and robustness, while underscoring the key role of numerical simulation in elucidating droplet-microstructure interactions.
keywords:
Pancake bouncing; SPH method; attractive forces; nonlocal contact repulsion
MSC Codes 37M05, 65M75, 70F10, 76M28.
@@ 1 Introduction
Understanding and controlling droplet rebound dynamics is essential for advancing liquid manipulation technologies in various engineering applications. This phenomenon is particularly critical in applications such as combustion systems [sirignano2014advances], inkjet printing [wijshoff2018drop], anti-icing surfaces [bird2013reducing], corrosion-resistant coatings [zhang2016superhydrophobic] and self-cleaning materials [blossey2003self]. Recent studies reveal that these technologies have broad applications across fields including aerospace, biomedicine, automotive engineering, and smart textiles [khan2022recent]. While hydrophobic coatings facilitate droplet removal, achieving directional control and efficient shedding requires not only low surface energy but also carefully designed nanostructures with tailored surface chemistry [nakajima2011design]. At such nano-enhanced interfaces, pancake bouncing [liu2014pancake] emerges as a remarkable phenomenon: droplets impacting nanostructured superhydrophobic surfaces lift off immediately in a flattened state without retracting, reducing contact time by approximately 75% compared to conventional rebound. Precise liquid control is crucial for advanced functional surfaces and devices, necessitating deeper research into droplet rebound dynamics, especially given their significant implications in healthcare and electronics.
@@
The dynamics of droplet rebound has garnered considerable scientific attention, driving a steady stream of key discoveries in interfacial science[mohammad2023physics]. Early studies by [schutzius2015spontaneous] revealed that droplets on superhydrophobic surfaces in low-pressure environments can self-remove through spontaneous levitation and trampoline-like bouncing. Later, [hao2015superhydrophobic] identified a superhydrophobic-like bouncing regime on thin liquid films, where the rebound dynamics were independent of the underlying substrate. Subsequent work by [liu2016superhydrophobic] introduced a facile method to fabricate hierarchical multilayered superhydrophobic surfaces for efficient droplet shedding. [yun2017bouncing] further explored symmetry-breaking impacts, showing that ellipsoidal droplets reduce contact time and suppress rebound magnitude. Recent advances have significantly enhanced control over droplet rebound behavior through novel interfacial strategies. [zhao2023golden] proposed a ’golden section’ design criterion to regulate rebound via structural spacing, challenging conventional approaches. [fang2023target] developed a leaf-inspired superhydrophobic cantilever for directional bouncing, while [lathia2023two] reduced contact time by coating droplets with hydrophobic particles. Additionally, through the utilization of innovative tools, a deeper understanding of the underlying mechanisms has been achieved. In their study, [kumar2024heat] utilized high-speed infrared thermography and a three-dimensional transient heat conduction COMSOL model to accurately map the dynamic heat flux distribution during droplet impact on a cold superhydrophobic surface. [goswami2023simultaneous] introduced a novel micro-controlled droplet generator capable of releasing two equally sized water droplets simultaneously on-demand, enabling the investigation of the behavior of two droplets on a dry substrate. [yue2024controllable] reported that controlled self-transport of bouncing droplets can be achieved with the assistance of ultra-smooth surfaces featuring wedge-shaped grooves. [qian2024emergence] uncovered molecular-scale wetting dynamics through in situ characterization, providing fresh insights into droplet-surface interactions. And [sobac2025small] conducted a comprehensive numerically assisted analysis based on verifiable assumptions such as quasi-stationarities and small Reynolds/Peclet numbers.
@@
Although experimental investigations provide critical observations, the numerical modeling is essential for achieving fundamental physical understanding and enabling technological applications. In fact, despite extensive experimental investigations into droplet impact phenomena, the fundamental physics governing collisions with hydrophobic surfaces remains inadequately understood, as noted in literature [khojasteh2016droplet]. On the contrary, numerical modeling offers unique insights into interfacial phenomena. For example, [huang2023phase] conducted numerical simulations based on the Cahn-Hilliard-Navier-Stokes model to achieve a complete and quantitative study on the effects of interfacial tension when it becomes sufficiently low. Additionally, [zhang2018numerical] simulated buoyancy-driven bubble-wall interactions using an adaptive ALE method, capturing the interaction of a rising bubble with a solid wall. Furthermore, [ray2024new] investigated the head-on collision of unequal-size droplets of the same liquid on wetting surfaces using direct numerical simulations at different size ratios. Based on the Onsager variational principle, [qian2024emergence] presented a phenomenological model that reveals how the intrinsic material parameters of soft gels dictate phase-separation dynamics. Moreover, fabricating nanostructured substrates for droplet impact studies incurs significant costs, whereas numerical simulations provide a cost-efficient alternative for parametric studies and experimental design optimization. Within this framework, numerical simulations have become pivotal in impact dynamics research. The creation of high-fidelity computational models capable of capturing intricate interfacial processes is essential for both theoretical advancements and practical applications.
@@
Smoothed Particle Hydrodynamics (SPH) method, as a mesh-free Lagrangian method, offers distinct advantages in simulating droplet rebound phenomena. In fact, droplet rebound typically involves large interfacial deformations and topological changes, presenting inherent difficulties for grid-based simulations. Traditional Eulerian approaches, such as the Lattice Boltzmann method (LBM), rely on multiphase domain discretization and phase-field method to accurately capture fluid interfaces. While LBM excels in modeling complex multi-phase flows, its application to gas-liquid systems with high density ratios demands non-trivial modifications to ensure numerical stability [QIAO20191216, qiao2024free, AAMM-17-4]. By contrast, SPH circumvents these limitations through its fully Lagrangian mesh-free formulation, eliminating the need for explicit gaseous-phase discretization altogether [king2023large]. This intrinsic capability renders SPH particularly suited for investigating droplet interactions with hydrophobic surfaces. Other alternative particle-based approaches such as molecular dynamics (MD) and dissipative particle dynamics (DPD) suffer from severe scalability limitations in mesoscale droplet simulations due to their demanding resolution requirements and consequent computational overhead. The SPH method achieves an optimal balance between numerical resolution and computational efficiency. Notably, while SPH has seen extensive practical applications in recent years, its theoretical foundations have been progressively strengthened. Significant contributions include Du et al.’s mathematical analysis from a nonlocal perspective [du2020mathematics, lee2019asymptotically] and Sun et al.’s stability analysis through an energy-based framework [feng2023energy, AAMM-14-5, zhu2024energy]. These theoretical advancements have collectively reinforced SPH’s position as a robust and well-validated methodology for droplet dynamics research.
@@
Surface tension modeling constitutes a critical component of SPH simulation for droplet dynamics, and numerous significant studies have been conducted in this area. Distinct from phase-field methods that regulate interface motion through energy minimization, SPH implementations necessitate the explicit definition of forces acting on interface particles to control interfacial behavior. In reality, researchers have developed two principal methodological approaches. The first approach is the Continuous Surface Force (CSF) formulation [brackbill1992continuum], which implements a volumetric force at fluid interfaces through local curvature estimation. While this method maintains physical consistency, it demands careful handling of surface normal vectors and curvature calculations near interface regions [huo2024modeling, dong2023simulation, blank2024surface]. Furthermore, this approach faces challenges in simulating droplet topological changes. The second approach involves the particle-particle interaction force (PIF) model [tartakovsky2005modeling, breinlinger2013surface], which bypasses the need for normal vector and curvature calculations. This method demonstrates versatility in simulating solid-liquid interactions and solid wettability. Despite its practical effectiveness, the second approach dependence on empirically determined strength of the interparticle attraction [dong2024droplet], lacking strong theoretical foundations, constitutes a major limitation. Notice that surface tension emerges from interparticle attractive forces. Thus, when these attraction strengths are empirically predetermined, the resulting surface tension coefficients become inherently uncontrollable, frequently leading to physical inconsistencies. Therefore, developing a method that preserves theoretical consistency while ensuring conciseness and broad applicability has become critically important.
@@
The contact algorithms is an another challenge for stable and precise SPH simulation, particularly for simulating droplet collisions on intricate arrays of microcones. The inherent complexity of droplet impact dynamics, characterized by multiscale solid-liquid interactions, creates escalating challenges for contact algorithm stability as the Weber number increases [pan2009binary]. The traditional repulsive force algorithms [huo2024modeling, dong2023simulation, dong2019modeling], as a popular contact algorithm, require accurate surface normals $\boldsymbol{n}$ and vertical distance $d_{p}$ to surface. Specifically, the repulsive force $\boldsymbol{F}_{i}^{\text{rep}}$ on particle $i$ satisfies the following relation:
(1) $\boldsymbol{F}_{i}^{\text{rep}}\propto\frac{2|d_{0}-d_{p}|}{(\Delta t)^{2}}m_{i}\boldsymbol{n}_{i},$
@@
where $\Delta t$ is time step, $m_{i}$ is particle mass and $d_{0}$ is contact threshold distance. Although this approach demonstrates satisfactory performance for ideally smooth substrates [mohammad2023physics], its extension to realistically rough substrates encounters fundamental limitations. A representative case involves micro-structured hydrophobic surfaces [liu2014pancake], as illustrated in Fig. 1.
@@
In fact, due to the non-smoothness of the substrates, it is hard to determinate the local surface normals $\boldsymbol{n}_{i}$ and vertical distance $d_{p}$. Besides, the virtual particle method is a widely used boundary treatment technique to address kernel truncation and wall penetration issues. However, when used in isolation, the virtual particle method often fails to reproduce physically accurate repulsive forces during droplet rebound. Consequently, the development of a robust contact algorithm capable of handling intricate microcone geometries is critical for reliable simulation of pancake bouncing dynamics.
@@ 2 Governing equations
(2) $\frac{\mathrm{d}\rho}{\mathrm{d}t}$ $=-\rho\nabla\cdot\boldsymbol{u},$
(3) $\rho\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}t}$ $=-\nabla p+\mu\nabla^{2}\boldsymbol{u}+\rho\boldsymbol{g}+\boldsymbol{f}_{\mathrm{att}}+\boldsymbol{F},$
The governing equations involve the fluid velocity field $\boldsymbol{u}$, hydrodynamic pressure $p$, and local densities $\rho$. The pressure $p$ is calculated by the following state of equation
(4) $p=\frac{c^{2}\rho_{0}}{\gamma}\left(\left(\frac{\rho}{\rho_{0}}\right)^{\gamma}-1\right).$
We adopt $\gamma=7$ [monaghan1994simulating] for the equation of state, with $\rho_{0}$ and $c$ denoting the reference densities and the numerical speed of sound, respectively. $\frac{\mathrm{d}}{\mathrm{d}t}$ represents the material derivative where $\frac{\mathrm{d}}{\mathrm{d}t}\equiv\frac{\partial}{\partial t}+\boldsymbol{u}\cdot\nabla$. $\mu$ denotes dynamic viscosity. The body forces consist of gravitation $\rho\boldsymbol{g}$, droplet cohesive attractive force $\boldsymbol{f}_{\mathrm{att}}$, and contact repulsion $\boldsymbol{F}$ at the boundary interface.
Accurate characterization of cohesive attractive and repulsive interactions is crucial for reliably simulating droplet rebound dynamics on microcone arrays surfaces. In this section, we present a unified mathematical framework that integrates smoothed kernel functions with nonlocal operators to describe both $\boldsymbol{f}_{\mathrm{att}}$ and $\boldsymbol{F}$, enabling robust three-dimensional modeling of bouncing behavior.
@@ 2.1 Modeling of droplet cohesion
It is well-established that the surface tension of water droplets originates from intermolecular forces (hydrogen bonding and van der Waals interactions) between water molecules. While intermolecular interactions are physically straightforward, they impose prohibitive computational costs in macroscopic system simulations. For surface tension modeling at continuum scales, cohesive force models offer a viable alternative. For the simplicity of the mathematical representation, we need to introduce the smoothing kernel function $W(\boldsymbol{x}-\boldsymbol{x}^{\prime},h)$, which satisfies the following conditions:
@@
  • Symmetric property: $W(\boldsymbol{x}-\boldsymbol{x}^{\prime},h)=W(\boldsymbol{x}^{\prime}-\boldsymbol{x},h)$;
  • Compact condition: $W(\boldsymbol{x}-\boldsymbol{x}^{\prime},h)=0\text{ where }|\boldsymbol{x}-\boldsymbol{x}^{\prime}|\geq h$.
Here, $|\cdot|$ represents the Euclidean norm and $h$ is the influence radius of kernel. $W(\boldsymbol{x},h)$ affects only a specific support domain. Additionally, the following lemma on quadruple integrals is essential for our subsequent analysis.
Lemma 2.1.
Assume $R_{0}>0$. Denote
$G[f]:=\int_{0}^{R_{0}}\int_{0}^{\pi}\int_{0}^{arc\cos(s/{R_{0}})}\int_{s/\cos(\theta)}^{R_{0}}f(r)\cos(\theta)r^{2}\sin(\theta)drd\theta d\phi ds.$
There holds
(5) $G[f]=\frac{\pi}{3}\int_{0}^{R_{0}}f(r)r^{3}dr.$
Furthermore, if $f(\|\mathbf{x}\|)=\partial_{\|\mathbf{x}\|}W(\mathbf{x},R_{0})$ and $W$ is any kernel function, there holds
$G[f]=\frac{1}{4}.$
The proof of Lemma 2.1 can be established using techniques of multiple integration. For completeness, we provide the detailed proof in the Appendix A.
We now proceed to construct a cohesive forces model capable of generating physically accurate interfacial tension. In our work, each Lagrangian particle represents a finite fluid volume element comprising numerous water molecules. Between adjacent particles, we introduce an artificial cohesive force term $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})$ to approximate the net effect of microscopic intermolecular forces. Specifically, we denote the attraction of the particle at $\boldsymbol{y}$ to the particle at $\boldsymbol{x}$ as follows
(6) $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})=A_{\sigma}\rho(\boldsymbol{y})\nabla_{\boldsymbol{x}}W(\boldsymbol{x}-\boldsymbol{y},{R_{\mathrm{att}}}),$
where $A_{\sigma}$ is a parameter dependent on interfacial tension $\sigma$ and $R_{\mathrm{att}}$ is the cohesive force interaction radius. The specific form of $A_{\sigma}$ will be given later. $\nabla_{\boldsymbol{x}}W(\boldsymbol{x}-\boldsymbol{y},{R_{\mathrm{att}}})$ in the same direction as $\boldsymbol{y}-\boldsymbol{x}$. Thus, by integrating over $\boldsymbol{y}$, the total attractive force acting on the particle located at position $\boldsymbol{x}$ is
(7) $\boldsymbol{f}_{\mathrm{att}}(\boldsymbol{x})=\int_{\Omega}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})d\boldsymbol{y}=A_{\sigma}\int_{\Omega}\rho(\boldsymbol{y})\nabla_{\boldsymbol{x}}W(\boldsymbol{x}-\boldsymbol{y},{R_{\mathrm{att}}})d\boldsymbol{y},$
@@
$V_{1}$$V_{2}$$\boldsymbol{x}_{0}$$\boldsymbol{t}$$R_{\mathrm{att}}$$R_{\mathrm{att}}$$L$